OpenStudy (anonymous):

simplify and express without radicals. Sqrt(x)/eleventh root (x) click for better format

OpenStudy (anonymous):

\[\sqrt{x}/\sqrt[11]{x}\]

OpenStudy (anonymous):

Rules: a^b * a^c = a^{b+c} a^b/a^c =a^{b-c} (a^b)^c = a^{bc} Numerator = sqrt x = x^(1/2) Denominator = x^(1/11) So, via the rules, x^(1/2-1/11) = x^(9/22) which u can "read" as the 22nd root of x raised to the power 9.